OpenStudy (anonymous):

Can anyone explain how to solve this, please? A publishing company has a weekly cost equation of C = 94,000 + 15x and a weekly revenue equation of R = 25x, where x is the number of books produced and sold in a week. The company loses money when R < C (the company breaks even when R = C). At what sales point does the company lose money?

OpenStudy (anonymous):

What does it mean if a business 'breaks even?'

OpenStudy (anonymous):

I don't really know.

OpenStudy (anonymous):

that's when the cost is equal to the income. They aren't making of losing any money. So I would find when that is and then say that anything less than that is losing money.

OpenStudy (anonymous):

Okay. But it doesn't really all make sense to me. Do I solve for x?

OpenStudy (anonymous):

@wilfley

OpenStudy (anonymous):

@texaschic101 ?

OpenStudy (texaschic101):

9400 + 15x = 25x 9400 = 25x - 15x 9400 = 10x 9400/10 = 940 they " break even " at 940 books so if there were 939 or less, then the company would lose money check... C side R side --------------------------- 9400 + 15x > 25x 9400 + 15(939) > 25(939) 9400 + 14085 > 23475 23485 > 23475 C > R

OpenStudy (anonymous):

It's 94,000

OpenStudy (texaschic101):

crap...hold on

OpenStudy (anonymous):

Is the answer the company loses money if it sells 9399 or fewer books?

OpenStudy (texaschic101):

94000 + 15x = 25x 94000 = 25x - 15x 94000 = 10x 94000/10 = x 9400 = x -- break even point if there were 9399 books... 94000 + 15x > 25x 94000 + 15(9399) > 25(9399) 94000 + 140985 > 234975 234985 > 234975 They would have to sell 9399 books (or less) to lose money

OpenStudy (texaschic101):

yes....company loses money if it sells 9399 books or less